OpenStudy (anonymous):

Use the elimination method to solve the system for x and y. 6x+2y=11 14x-3=y Select one: a. (1/2, 4) b. (4, 2) c. (4, 1/2) d. (2, 4)

OpenStudy (anonymous):

6x+2y=11 14x-3=y (rearrange and multiply by 2) therefore 6x+2y=11 + 28x-2y=6 ________________ 34x =17 x=17/34 x=1/2 substitute x=1/2 into 1: 6(1/2)+2y=11 3+2y=11 2y=11-3 2y=8 y=8/2 y=4 Therefore answer is a. (1/2,4)

OpenStudy (anonymous):

14x-3=y rearrange make 14x-y=3 then multipy by 2: 28x-2y=6